Colloquium

A Mathematics Inquiry Learning at Kutztown (MILK) Lecture

5:00 p.m. Thursday, October 14, 2010

Boehm Hall 261

An Invitation to a Journey in Inquiry Based Learning

Dr. Stan Yoshinobu California Polytechnic University at San Luis Obisbo Associate Professor of Mathematics

Abstract

At the end of the day, what ultimately matters in a class, any class, is what students do. Inquiry-based learning (IBL) is a method of teaching Mathematics that is as its core learner centered. In an IBL classroom, students activity construct mathematical ideas and arguments. Students learn more than just factual knowledge. They also develop the ability to think critically, to explore, to discover, to reflect, to be sociable and effectively communicate, to get unstuck, and all those other wonderful traits that we value.

This talk will be divided into two parts. In the first part, I will highlight the scientific reasons why we should use IBL. In the second part, videos of an actual IBL course will be used to generate a discussion about what an actual IBL course looks like.

4:30 p.m. refreshments served

5:00 p.m. talk begins

The Mathematics Inquiry Learning at Kutztown (MILK) Lecture is funded by the generous support of the Academy for Inquiry-Based Learning, The Educational Advancement Foundation, and Mr. Harry Lucas, Jr.

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache III, Chair